It is possible for the exterior angles of a regular polygon to measure 30 degrees.
True or False?
It is possible for the interior angles of a regular polygon to measure 145 degrees.
True or False?
Calculate the measure for each exterior angle of a regular nonagon.
Answers: 20 degrees, 40 degrees, 140 degrees, 1260 degrees.
calculate the measure for each interior angle of a regular 15-gon.
Ansers: 12 degrees, 24 degrees, 156 degrees, 2340 degrees.

Answer by stanbon(48535)

It is possible for the exterior angles of a regular polygon to measure 30 degrees.
True or False?
True if the polygon has 12 sides.
------------------------------------------------
It is possible for the interior angles of a regular polygon to measure 145 degrees.
True or False?
False because 360/145 is not a whole number.
-------------------------------------------------
Calculate the measure for each exterior angle of a regular nonagon.
360/9 = 40 degrees
Answers: 20 degrees, 40 degrees, 140 degrees, 1260 degrees.
-------------------------------------------------------
calculate the measure for each interior angle of a regular 15-gon.
Each EXTERIOR angle is 360/15 = 24 degrees
So each INTERIOR angle is 180-24 = 156 degrees
Ansers: 12 degrees, 24 degrees, 156 degrees, 2340 degrees.
=================================================================
Cheers,
Stan H.